OpenStudy (anonymous):

Given the function g(x) = −2x + 4, find g(−5). −6 4.5 6 14

OpenStudy (anonymous):

i think its -6 but im not sure

OpenStudy (anonymous):

It would be 14. -2(-5)+ 4= 10+4= 14

OpenStudy (anonymous):

oh ok that makes more sense haha